OpenStudy (anonymous):

An aircraft has a mass of 3.60 x 105 kg. At a certain instant during its landing, its speed is 50.6 m/s. If the braking force is 3.20 x 105 N, what is the speed of the airplane 19.4 s later?

OpenStudy (anonymous):

Is the braking 3.2x10^5N?

OpenStudy (anonymous):

Anyway, Equation: F=ma You need acceleration in order to calculate velocity. So we get: 3.2x10^5N=3.6x10^5*a <-- isolate a a= 0.88 m/s^2 (This will BE NEGATIVE BECAUSE BRAKE FORCE IS WORKING AGAINST US) v = u +at (Let u be initial velocity and t (be time) v= 50.6+(-0.88)(19.4)=32.98 m/s Hope this helps :)